# ATNPA: A Unified View of Oversmoothing Alleviation in Graph Neural Networks

### v1

@@ Abstract
Oversmoothing is a commonly observed challenge in graph neural network (GNN) learning, where, as layers increase, embedding features learned from GNNs quickly become similar/indistinguishable, making them incapable of differentiating network proximity. A GNN with shallow layer architectures can only learn short-term relation or localized structure information, limiting its power of learning long-term connection, evidenced by their inferior learning performance on heterophilous graphs. Tackling oversmoothing is crucial to harness deep-layer architectures for GNNs. To date, many methods have been proposed to alleviate oversmoothing. The vast difference behind their design principles, combined with graph complications, make it difficult to understand and even compare their difference in tackling the oversmoothing. In this paper, we propose ATNPA, a unified view with five key steps: Augmentation, Transformation, Normalization, Propagation, and Aggregation, to summarize GNN oversmoothing alleviation approaches. We first outline three themes to tackle oversmoothing, and then separate all methods into six categories, followed by detailed reviews of representative methods, including their relation to the ATNPA, and discussion about their niche, strength, and weakness. The review not only draws in-depth understanding of existing methods in the field, but also shows a clear road map for future study.
@@ 1 Introduction
Graph Neural Networks (GNN) [1, 2, 3] have become prevalent in support learning from networked data, especially after the success of the Graph Convolution Network (GCN) [4]. The main goal of GNN is to learn feature representation [5] for network entities, such as nodes or edges, in order to support downstream tasks, like node classification or link prediction. While GNN has achieved competitive performance in many benchmark graph datasets, it is known to only perform well with shallow layer architectures and cannot learn long-term node-node relation well. One consequence of such inability leads to its inferior performance on heterophilous graph [6].
It has been shown that simply stacking GNN layers to build a deep architecture cannot learn well due to the observed oversmoothing phenomenon [7, 8]. GNN models equipped with oversmoothing alleviation can in general accommodate more GNN layers and therefore allow nodes to have a larger receptive fields [9]. As a result, models aiming to alleviate oversmoothing also tend to gain advantage over heterophilous datasets. Such dual relation has been observed in several studies [10, 11].
@@
Oversmoothing can be described as a phenomenon that all node mebeddings, after deep GNN layers, become similar to each other. Several measures such as Dirichlet energy [12] and Mean Average Distances (MAD) have been proposed to quantify the extent of oversmoothing of a model [13]. In this paper, unless otherwise specified, Dirichlet energy will be used as the major measure of oversmoothing for analysis.
Majority theoretical works studying when oversmoothing occurs are based on dynamic systems. One important work [14] uses subspace theorem to analyze when GCN becomes oversmoothing asymptotically (i.e., infinite layers behavior). This theory motivates many alleviation approaches, such as constrain-based and random edge dropping such as EGNN [15] and edge-drop [16]. Another work [12] uses perturbation theory to analyze how Dirichlet energy (energy of the dynamic system) behaves under different coupled systems [12]. This theory mostly serves as basis for dynamic-based alleviation approaches, such as graphCON [12] and G2-gating [11]. Recently, another theoretical work [17] suggests that graph attention network (GAT) [18] also suffers from oversmoothing. The first theory guides many constrain-based work and random edge dropping such as EGNN [15], edge-drop [16] while the second work mostly serves dynamic-based models such as graphCON, G2-gating [11].
@@
Indeed, many works have been proposed to tackle oversmothing in GNNs, by using different types of design principles. For example, energy based approaches aim to increase initial energy or keep energy from exponential decay during propagation in GNNs. On the other hand, other methods focus on decoupling topology propagation and feature transformation. The vast difference in their design principles, combined with complicated graph topology and message passing mechanisms, making it difficult to understand and even compare their difference in tackling the oversmoothing. Very few survey papers exist to review methods tackling GNN oversmoothing challenges. A recent survey [19] has compared several methods, and pointed out that some of the existing methods (such as GCNII [20], GraphCon [12]) cannot increase their model performance with deep layers despite the oversmoothing measure (i.e., Dirichlet energy) is preserved to be constant among layers, mainly because of lacking expressive power. Therefore, existing survey [19] is mainly focused on reviewing method drawback from the expressive power perspective.
@@
In this paper, we propose to unify existing methods to the same form and point out their potential connection in tackling GNN oversmoothing. Our study not only provides a unified view, ATNPA  to summarize all methods using common math formulations, but also separate them into six groups, by taking their unique designs into consideration. The survey outlines differences between methods in each groups, explains their rationality, and addresses their limitations. Our review has a number of math formulas, because reviewed papers are heavy in math formulations. To precisely summarize and highlight their difference, we keep representative methods’ backbone formulas in the review for a better understanding.
@@ 2 Problem Notation
$H^{l}\leftarrow A^{l-1}\texttt{f}_{\theta}(X);$ (1)
@@
Where Eq. (1) can be considered as a linear Graph Convolution as proposed by [21] and Eq. (2) is a common graph convolution propagation.
@@
We use $\texttt{LA}(\cdot)$ to denote layer-wise aggregations that aggregate embeddings learnt from current and preceding layers. Examples of aggregation include concatenate, max pooling, and LSTM-attention operations [25]. We use $\cup$ to denote the input is a union set of embeddings from all layers.
@@
A common choice of $\texttt{Aug}(\cdot)$ could be symmetric Laplacian, Laplacian, First-order Chebshev approximation (akin to GCN) following traditional spectral graph theory. Other choices include different random masking schemes such as random edge dropping [16] which is proven to be effective both empirically [16] and theoretically [14], and learnable attention matrix that has the same structure as A (examples include transformer architecture [26] and diffusivity in GRAND [27]).
@@ 2.1 Oversmoothing Definition
Oversmoothing indicates that $D_{M}(H^{l})\to 0$ exponentially converges, w.r.t the increase of layer value $l$.
@@
The definition above is similar to the definition in [14] with $\mu(\cdot)$ defined as $D_{M}(\cdot)$. The second definition is often used in diffusion-based system analysis such as [12] with $\mu$ as the Dirichlet energy of the system while the first definition is often used in GNN-backbone methods such as EGNN [15] and DropEdge [16].
@@ 2.2 Oversmoothing Measures
Note that MAD is closely related to cosine similarity and therefore only considers the direction of the two embeddings and ignores their magnitude difference. Compared with Dirichlet energy measure, caution on feature magnitude is needed when using the MAD measure.
2.3 Model Expressive Power
Model expressive power and oversmoothing have intrinsic connections. Under this assumption, models with strong express power may be more resilient to oversmoothing. Therefore, enhancing model expressive power seems to be also helpful for oversmoothing alleviation.
In the graph neural network domain, existing research on model expressive power is primarily focused on comparing GNNs’ separation ability with Weisfeiler-Lehman (WL) graph isomorphism test [28]. Theoretical analysis for existing studies is built upon two important concepts: (1) multi-set; or (2) permutation invariant injective aggregation function [29] [30]. For the former, the neighbor set along with the center node is formed as a multi-set, on which the expressive power is then defined accordingly. For the latter, aggregation function should be injective so that non-isomorphic graphs can be mapped to unique embeddings and permutation invariant property guarantees that isomorphic graphs (which can be obtained by permutation group action) can be mapped to embeddings belonging to the same equivalent class.
While graph isomorphism test focuses on graph-level classification, we can consider node-level classification as a local subgraph differentiation test. For a graph $G_{i}$, we use $\mathbb{P}(G_{i})$ to denote the set of all graphs isomorphic to $G_{i}$. For one $L$ layer GNN $F_{\theta}(\cdot):G\rightarrow R^{d}$, if two graphs $G_{1}$, $G_{2}$ expanded from two target nodes (i.e. each graph includes target node and its $L$ hop neighbors) are non-isomorphic, we can define an $\varepsilon_{\text{MEP}}$ score in Eq. (10) to quantify model expressive power over node separation ability:
$\varepsilon_{\text{base}}(G_{i})$ $=\sup_{G\in\mathbb{P}(G_{i})}||F_{\theta}(G_{i})-F_{\theta}(G)||_{2}^{2}$ (9)
$\varepsilon_{\text{MEP}}$ $=\frac{||F_{\theta}(G_{1})-F_{\theta}(G_{2})||^{2}_{2}}{\max(\varepsilon_{\text{base}}(G_{1}),\varepsilon_{\text{base}}(G_{2}))+\epsilon}$ (10)
where $\epsilon$ is a scalar ensuring that the denominator does not equal to zero, which happens when $F_{\theta}(\cdot)$ is permutation invariant.
Overall, Eq. (10) defines that for non-isomorphic graphs $G_{1}$ and $G_{2}$, the larger the $\varepsilon_{\text{MEP}}$ score of the model, the better its expressive power is in separating $G_{1}$ and $G_{2}$. Meanwhile, because $F_{\theta}(G_{1})$ denotes embedding of graph $G_{1}$, oversmoothing would imply that $F_{\theta}(G_{1})\approx F_{\theta}(G_{2})$, resulting in a small $\varepsilon_{\text{MEP}}$ score for non-isomorphic graphs.
Assuming a model has an infinitely strong power of differentiating node up to $L$-hops away, it implies that the model can avoid oversmoothing up to $L$-layers. This seems to suggest that oversmoothing is a side effect brought by the model lacking expressive power in separating nodes (including node and its local neighborhood). A recent study [31] provides a positive evidence towards this hypothesis by showing that increasing the node expressive power through local structure aware GNN can help alleviate oversmoothing issue.
@@ 2.3 Model Expressive Power
Methods Category Energy (Rewiring) Energy (Nomalization) Energy (Coefficient) Energy (Initialization) Decoupling Dynamics Dense/Residual
ResGCN [32] residual-based
APPNP [33] residual-based
GCNII [20] residual-based
@@
EGNN [15] residual-/energy-based
GroupNorm [35] residual-/energy-based
G2-gating [11] residual-/energy-/diffusion-based
@@
DAGNN [36] dense-based
DCGCN [37] dense-based
MixHop [38] dense-based
@@
DropConnect [39] random-mask based
DropMessage [40] residual-/random-mask based
@@
NodeNorm [24] energy-based
@@
GraphCon [12] diffusion-based
ACMP  [41] diffusion-based
Neural Sheaf Diffusion [42] diffusion-based
GraphT [26] transformer-based
GraphiT [43] transformer-based
3 ATNPA: A Unified View of Oversmoothing Alleviation
In this session, we first outline message propagation process commonly used in GNN learning (Sec 3.1), then summarize main themes to tackle oversmoothing (Sec 3.2), which help lay the foundation for ATNPA, the unified view and categorization (Sec 3.3). Secs 3.4 and beyond review representative methods in each category, including their key steps and relation to the ATNPA, as well as their rationality in tackling the oversmoothing challenges.

### v2

@@ Abstract
Oversmoothing is a common challenge in learning graph neural networks (GNN), where, as layers increase, embedding features learned from GNNs quickly become similar or indistinguishable, making them incapable of differentiating network proximity. A GNN with shallow layer architectures can only learn short-term relation or localized structure information, limiting its power of learning long-term connection, evidenced by their inferior learning performance on heterophilous graphs. Tackling oversmoothing is crucial for harnessing deep-layer architectures for GNNs. To date, many methods have been proposed to alleviate oversmoothing. The vast difference behind their design principles, combined with graph complications, make it difficult to understand and even compare the difference between different approaches in tackling the oversmoothing. In this paper, we propose ATNPA, a unified view with five key steps: Augmentation, Transformation, Normalization, Propagation, and Aggregation, to summarize GNN oversmoothing alleviation approaches. We first propose a taxonomy for GNN oversmoothing alleviation which includes three themes to tackle oversmoothing. After that, we separate all methods into six categories, followed by detailed reviews of representative methods, including their relation to ATNPA, and discussion of their niche, strength, and weakness. The review not only draws an in-depth understanding of existing methods in the field but also shows a clear road map for future study.
keywords:
Oversmoothing, graph neural networks, graph embedding, graph learning, review
@@ 1 Introduction
Graph Neural Networks (GNN) [1, 2, 3] have become prevalent in support learning from networked data, especially after the success of the Graph Convolution Network (GCN) [4]. The main goal of GNN is to learn feature representation [5] for network entities, such as nodes or edges, in order to support downstream tasks, like node classification or link prediction. While GNN has achieved competitive performance in many benchmark graph datasets, it is known to only perform well with shallow layer architectures but cannot learn long-term node-node relation well. One consequence of such inability leads to its inferior performance on heterophilous graph [6].
@@
It has been shown that simply stacking GNN layers to build a deep architecture cannot learn well due to the observed oversmoothing phenomenon [7, 8]. Oversmoothing can be described as a phenomenon that all node emebeddings, after deep GNN layers, become similar to each other. Several measures such as Dirichlet energy [9] and Mean Average Distances (MAD) have been proposed to quantify the extent of oversmoothing of a model [10]. In this paper, unless otherwise specified, Dirichlet energy will be used as the major measure of oversmoothing for analysis. Figure 1 demonstrates the GCN embedding learning results from the Karate network [11], where increasing GCN layers from 1 to 3 results in better class separability whereas increasing the layer further results in embedding features with deteriorated class separability. Figure 2(b) demonstrates the oversmoothing phenomenon on the Cora network [12] where all node embedding becomes similar to each other.
GNN models equipped with oversmoothing alleviation can in general accommodate more GNN layers and therefore allow nodes to have a larger receptive fields [13]. As a result, models aiming to alleviate oversmoothing also tend to gain advantage over heterophilous datasets. Such dual relation has been observed in several studies [14, 15].
1.1 Research Gap and Motivations
@@
Indeed, many works have been proposed to tackle oversmoothing in GNNs, by using different types of design principles. For example, energy-control approaches aim to increase initial energy or keep energy from exponential decay during propagation in GNNs. Other methods focus on decoupling topology propagation and feature transformation. The vast difference in their design principles, combined with complicated graph topology and message passing mechanisms, making it difficult to understand and even compare their difference in tackling the oversmoothing. Very few survey papers exist to review methods tackling GNN oversmoothing challenges. A recent survey [16] has compared several methods, and pointed out that some of the existing methods (such as GCNII [17], GraphCon [9]) cannot increase their model performance with deep layers despite the oversmoothing measure (i.e., Dirichlet energy) is preserved to be constant among layers, mainly because of lacking expressive power. Therefore, existing survey [16] is mainly focused on reviewing method drawback from the expressive power perspective.
@@
Classical GNN message passing includes two critical operations: message aggregation and message update, where message aggregation focus on how to aggregate information passed from connected neighbors and message update focuses on how to update the node information to the next iteration or next layer. In many works, augmentation prior to training, such as feature or edge dropout, is also a part of standard pipeline, which can be simply described as a stochastic masking over original data sources, including graph topology and node features for graph learning.
To alleviate oversmoothing for GNN learning, most existing methods are motivated by modifying or changing the three operations, including augmentation, message aggregation, and message update, in the GNN training process with the objective of obtaining final embeddings without oversmoothing. Nevertheless, there is no systematic study of existing methods on how the three main operations affect the oversmoothing problem. To close the gap and provide guidance for future studies of oversmoothing, we propose a taxonomy with a unified view of five key steps: Augmentation, Transformation, Normalization, Propagation, and Aggregation, to summarize GNN oversmoothing alleviation approaches. The unified view therefore provides a systematic study of the three operations (augmentation, message aggregation, and message update) with more detailed analysis and categorization of existing methods.
We notice that the three major themes only correspond to three main operations in classical GNN but also reflect the underlying different principles that are leveraged to alleviate oversmoothing. The methods we selected, despite various motivations, inspirations, and structures, all fall into the three major themes and their underlying principles.
1.2 Contributions
@@
In this paper, we propose to unify existing methods in the same form and study their connections in tackling GNN oversmoothing. Our study not only provides a taxonomy and a unified view, ATNPA  to summarize all methods using common math formulations, but also separates them into three themes and six subgroups, by taking their unique designs into consideration. The survey outlines the differences between methods in each group, explains their rationality, and addresses their limitations. Our review has a number of math formulas, because reviewed papers are heavy in math formulations. To precisely summarize and highlight their difference, we keep representative methods’ backbone formulas in the review for a better understanding.
The remainder of the paper is structured as follows. Section 2 provides the definition of oversmoothing problem and its common measures in existing works. Section 3 describes our taxonomy for GNN oversmoothing problems from different angles. Section 4 introduces our main unified view and the six categories of existing methods, followed by a detailed discussion over each category. Section 5 concludes our works along with future guidelines for oversmoothing problems. For ease of reference, Table 1 summarizes key symbols and notations used in the paper.
(a) Karate Club Topology
(b) 1-layer GCN Embedding
(c) 3-layers GCN Embedding
(d) 5-layers GCN Embedding
(e) 10-layers GCN Embedding
(f) 20-layers GCN Embedding
Figure 1: Visualization of GCN node embedding results with respect to increasing number of layers (Karate club network [11]). Nodes are color-coded with same colored nodes belonging to the same class. From top to bottom, left to right, (a) denotes the Karate club network toppoloy, (b) to (f) denotes node embedding from 1-layer GCN to 20-layer GCN, respectively. As GCN layer increases from 1-layer to 3-layer, the embedding achieve better class separability (i.e., better results). As layer continuously, from 5-layers to 20-layers, GCN embedding loss node separability.
(a) 1-layer GCN Embedding
(b) 3-layers GCN Embedding
(c) 5-layers GCN Embedding
(d) 10-layers GCN Embedding
(e) 15-layers GCN Embedding
(f) 20-layers GCN Embedding
Figure 2: Visualization of oversmoothing from GCN embedding learning (Cora network [12]). Nodes are color-coded with same colored nodes belonging to the same class. From top to bottom, left to right, (a) to (f) denotes node embedding from 1-layer GCN to 20-layer GCN, respectively. As GCN layer increases from 1-layer to 5-layer, the embedding achieve better class separability (i.e., better results). As layer continuously, from 10-layers to 20-layers, GCN embedding loss node separability.
Table 1: Summary of key symbols and notations.
Notations Descriptions
$G=\left(V,E,X\right)$ An attributed graph with node set ($V$), edge set ($E$), node feature content ($X$)
$A\in\mathbb{R}^{n\times n}$ An adjacency matrix
$\tilde{A}$ An augmented or synthesized adjacency matrix
$\sigma(\cdot)$ A non-linearity activation function
$H$, $H^{\prime}$, $H^{\prime\prime}$ Graph node embedding, the changing rate of $H$ (first derivative), and the changing rate of $H^{\prime}$ (second derivative)
$H^{l}\in\mathbb{R}^{n\times f}$ Graph node feature embedding learned at the $l^{th}$ layer
$H_{c}^{l}$ Graph node feature embedding learned through convolution (or in general any local aggregation)
$U$ $U$ denotes the velocity of embedding changing (assume graph node feature embedding evolves continuously with time $t$)
$\texttt{f}_{\theta}(\cdot)$ A content-based feature encoder parameterized by $\theta$
$\texttt{F}_{\theta}(\cdot)$ A graph convolution operator parameterized by $\theta$
$\texttt{NT}(\cdot)$ Normalization techniques in GNNs
$\texttt{LA}(\cdot)$ Layer-wise aggregations
$\texttt{Aug}(\cdot)$ Topology augmentation function
@@ 2 Problem Notation
$H^{l-1}\leftarrow\texttt{f}_{\theta}(H^{l-1});$ (1)
@@
Where Eq. (1) represents encoding node feature only before propagation and Eq. (2) shows the propagation process after feature encoding.
@@
We use $\texttt{LA}(\cdot)$ to denote layer-wise aggregations that aggregate embeddings learnt from current and preceding layers. Examples of aggregation include concatenate, max pooling, and LSTM-attention operations [22].
@@
Notice that Eq. (4) often occurs in dense-based approach and can be treated as an ensemble trick over different layers.
@@
A common choice of $\texttt{Aug}(\cdot)$ could be symmetric Laplacian, Laplacian, First-order Chebyshev approximation (akin to GCN) following traditional spectral graph theory. Other choices include different random masking schemes such as random edge dropping [23] which is proven to be effective both empirically [23] and theoretically [19], and learnable attention matrix that has the same structure as A (examples include transformer architecture [24] and diffusivity in GRAND [25]).
@@ 2.1 Oversmoothing Definition
where $\|\cdot\|^{2}_{F}$ is the matrix norm. Oversmoothing indicates that $D_{M}(H^{l})\to 0$ exponentially converges, w.r.t the increase of layer value $l$.
@@
The definition above is similar to the definition in [19] with $\mu(\cdot)$ defined as $D_{M}(\cdot)$ with the measure decay rate limit to exponential decay. The second definition is often used in diffusion-based system analysis such as [9] with $\mu$ as the Dirichlet energy of the system while the first definition is often used in GNN-backbone methods such as EGNN [26] and DropEdge [23].
@@ 2.2 Oversmoothing Measures
where $H_{i}^{l},H_{j}^{l}$ are feature vectors reflecting nodes $L_{2}$ distance with respect to its neighbors. In practice, the specific distance metric can be treated as a learnable parameter or hyperparameter.
@@
where we can observe that MAD simply replaced $L_{2}$ distance metric to cosine similarity compared with Eq. (7). Note that MAD is closely related to cosine similarity and therefore only considers the direction of the two embeddings and ignores their magnitude difference. Compared with Dirichlet energy measure, caution on feature magnitude is needed when using the MAD measure.
Table 2: Common dataset used for existing methods. Only datasets that are used at least twice are collected. Year reflect the time that the method is proposed. # of Total Usage indicates the number of times the dataset is used for existing method.
Methods/Datasets Year Cora [27] Citeseer [27] Pubmed [27] Texas [28] Cornell [28] Wisconsin [28] CoauthorCS [29] Coauthors-Physics [29] Chameleon [30] Amazon Photos [29] Squirrel [30] ogbn-arxiv [31] [32] Amazon Computers [29] Film [28] Cora-ML [33] Reddit [34]
APPNP 2018 $\checkmark$ $\checkmark$ $\checkmark$
GCNII 2020 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
GEN 2020 $\checkmark$
GroupNorm 2020 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
EGNN 2021 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
G2-gating 2023 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
GRN 2024 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
JKnet 2018 $\checkmark$ $\checkmark$ $\checkmark$
MixHop 2019 $\checkmark$ $\checkmark$ $\checkmark$
DAGNN 2020 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
DropEdge 2020 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
DropConnect 2020 $\checkmark$ $\checkmark$ $\checkmark$
DropMessage 2023 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
PairNorm 2020 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
NodeNorm 2020 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
WeightRep 2024 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
GRAND 2021 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Neural Sheaf Diffusion 2022 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
GraphCon 2022 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACMP 2023 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
# of Total Usage \ 17 17 16 7 7 7 5 5 5 5 4 4 4 3 2 2
# of Nodes \ 2708 3327 19717 183 183 251 18333 34493 2277 7650 5201 169343 13752 7600 2810 232965
2.3 Oversmoothing Benchmarks
Table 2 lists commonly used benchmark datasets by existing methods for oversmoothing validation. The datasets are listed and ordered, from left to right, based on the number of times each dataset is used in descending order. Cora, Citeseer, and Pubmed datasets are three common homophilic benchmarks used. Texas, Cornell, Wisconsin are three common heterophilic benchmarks used. Frequently used datasets are all small or medium scale ranging from $10^{3}$ to $10^{4}$ number of nodes. Few methods test shared large-scale heterophilic or homophilic benchmarks over $10^{5}$ number of nodes.
Oversmoothing problem occurs when GNN layers are stacked, aiming to learning global and longer hop-size node relations. For node classification tasks, homophilic graphs assume that nodes intend to share same labels with their nearby neighbors and the heterophilic graphs have more nodes sharing same labels with distant hops away. Heterophilic graphs could show more benefits and improvements for the models with deep GNN layers and reflect the model’s ability to learn the global node relationship[15]. Therefore, we recommend that future studies focus more on datasets that are heterophilic for testing the effectiveness of oversmoothing alleviation. Moreover, it can be observed that few large-scale heterophilic datasets are commonly tested by existing methods, resulting a lack of recognition of scalability comparison among existing methods.
@@ 2.3 Oversmoothing Benchmarks
Methods Category Energy(Rewiring) Energy(Nomalization) Energy(Coefficient) Energy(Initialization) Decoupling Dynamics Residual Dense
ResGCN [35] residual connection
APPNP [36] residual connection
GCNII [17] residual connection
@@
EGNN [26] residual connection/energy control
GroupNorm [38] residual connection/energy control
GRN [39] residual connection/energy control
G2-gating [15] residual connection/energy control/physics-inform
@@
DAGNN [40] dense concatenation
DCGNN [41] dense concatenation
MixHop [42] dense concatenation
@@
DropConnect [43] stochastic masking
DropMessage [44] residual connection/stochastic masking
@@
NodeNorm [21] energy control
WeightRep [45] energy control
@@
GraphCon [9] physics-inform
ACMP [46] physics-inform
Neural Sheaf Diffusion [47] physics-inform
GraphT[24]/GraphiT [48] graph transformer
3 A Taxonomy for GNN Oversmoothing Alleviation
In this section, we first outline message propagation process commonly used in GNN learning (Sec 3.1), then summarize main themes to tackle oversmoothing (Sec 3.2). After that, we propose a taxonomy for GNN oversmoothing alleviation, as shown in Fig. 3.